Circular No. 6968

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cv IN ESO 237-G42 J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova (B ∼ 17.5) by Roberto Antezana on a T-Max 400 film taken by L. E. González using the Maksutov telescope at Cerro El Roble on June 24.315 UT. The supernova is located at α = 22^{h}09^{m}46^{s}.29, δ = −49^{o}47′43′′.0 (equinox 2000.0), which is 5′′.0 east and 12′′.7 north of the nucleus of the spiral galaxy ESO 237-G42 (IRAS 22064–5003). SN 1998cv is not present on a film obtained with the same telescope on May 30.355 (B > 20). The object was confirmed spectroscopically by M. Phillips at Cerro Tololo using the Blanco 4-m telescope on June 27; he reports that this appears to be a type-Ic supernova relatively near maximum. The spectrum also shows narrow Hα emission at a redshift (∼ 8150 km/s) consistent with that of the host galaxy.

SUPERNOVA 1998cw IN ANONYMOUS GALAXY Maza also reports that another supernova (B ∼ 17.5) was found by

Antezana on a film taken as above on June 24.365 UT. SN 1998cw is located at α = 22^{h}56^{m}19^{s}.67, δ = −35^{o}09′59′′.9 (equinox 2000.0), which is on top of a small, faint galaxy (which is smaller than 3′′ on the Digitized Sky Survey). SN 1998cw is not present on a film obtained with the same telescope on May 30.374 (B > 20), but it was confirmed by Phillips with the the Blanco 4-m telescope on June 27. Phillips notes: “This is a type-Ia supernova, probably caught a few days before maximum. The spectrum resembles the pre-maximum spectrum of SN 1991T in showing strong Fe III absorption lines. However, the Ca II H and K and Si II 635.5-nm lines are also strong. A 20-min spectrum of the faint galaxy near this supernova does not show any obvious emission or absorption features.”

CH CYGNI J. Mikolajewska, Copernicus Astrnomical Center, Warsaw; R. J. Ivi-

son, Institute for Astronomy, Edinburgh; and R. Neri, IRAM, St. Martin d’Heres, write: “Continuum measurements with the IRAM interferometer and with SCUBA on the James Clerk Maxwell Telescope suggest that the symbiotic binary CH Cyg (cf. IAUC 6967) has undergone a new radio outburst. In May 1997 and Jan. 1998, the 1–3-mm spectrum was flat, with flux densities of 26 ± 3 at 1 mm and 23 ± 2 mJy at 3 mm, while on May 18 we observed a steeply rising continuum with flux densities of 39 ± 6 at 2 mm and 90 ± 7 mJy at 0.85 mm.”

1998 July 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green